# Reducing GPU Memory Fragmentation via Spatio-Temporal Planning for Efficient Large-Scale Model Training

### v1

@@ 8.4. Performance Breakdown
To understand the performance contribution of the static and dynamic allocators in STWeaver, we evaluate the performance breakdown of STWeaver when training the Qwen1.5-MoE-A2.7B model with the same setting in §8.2. To this end, we sequentially disable the dynamic allocator reusing Static Allocation Plan (mentioned as STWeaver w/o reuse), and the static allocator (mentioned as Caching Allocator, which is the vanilla PyTorch caching Allocator), and measure the corresponding memory efficiency in the above cases.
Static Allocator. The results in Figure 12 indicate that STWeaver with only the Static Allocation Plan reduces fragmentation memory by 70.2% compared to PyTorch Caching Allocator. This reduction in fragmentation memory accounts for 91% of the total fragmentation memory reduction achieved by the complete STWeaver system relative to PyTorch. Static planning accounts for the predominant share of the defragmentation result, primarily because static memory allocations form a substantial majority (from 73.4% to 99.3%) of the total memory allocation, as shown in Table 3.
@@
Dynamic fallback w/o reuse (GB) 15.19 1.61 17.80 1.78 2.97 1.95
Dynamic fallback with reuse (GB) 15.19 1.12 17.22 1.70 1.65 1.55
Dynamic Allocator. Compared with STWeaver without dynamic reuse, the full STWeaver reduces memory fragmentation by an additional 22.9%, mainly by lowering fallback allocations to the caching allocator. As shown in Table 3, enabling dynamic reuse decreases the number of requests falling back to the caching allocator. This benefit is most evident under recomputation, where caching allocations drop by 24.9%. Without recomputation, the impact is smaller.
@@ 9. Related Work
Online GPU Allocators. To reduce memory fragmentation and improve allocation efficiency, a plethora of online GPU memory allocators (Adinetz and Pleiter, 2014; Widmer et al., 2013; Steinberger et al., 2012; Vinkler et al., 2014) have been developed. Dynamic allocators operate atop the native GPU memory APIs (e.g., cudaMalloc) in a similar manner as the caching allocator of PyTorch. Differently, such allocators are meant to run on GPU threads alongside GPGPU applications, rather than managing GPU memory from the host like PyTorch and GMLake. To reduce fragmentation, they usually adopt sophisticated allocation policies such as the slab and buddy systems. Also, to achieve high-throughput allocations on GPUs, they have proposed scalable synchronization primitives across the massive threads of GPUs (Gelado and Garland, 2019). As a pluggable allocator of PyTorch, STWeaver also chooses to manage GPU memory from the host to improve usability and programmability.
Generic Memory Defragmentation Techniques. Memory defragmentation has been studied and discussed in various scenarios (Aigner et al., 2015; Johnstone and Wilson, 1998; Krauss et al., 2023) beyond GPU applications. Previous work (Veldema and Philippsen, 2012; Marlow et al., 2008; Hudson and Moss, 2001; Siegwart and Hirzel, 2006) has proposed defragmentation strategy based on data movement or copying. These approaches are mainly deployed in the real-time system with unpredictable runtime behaviors, which result in complex defragmentation strategies with high runtime overhead.
Machine Learning Compilers. Machine Learning compilers that convert high-level computation graphs to GPU instructions must manage memory allocation for these graphs. The compilers pre-analyze control dependencies to optimize memory layouts before execution. Current compilers like TVM (Chen et al., 2018) and TFLite (David et al., 2021) use greedy heuristics for reasonable allocation, while Checkmate (Jain et al., 2020) employs solvers for optimal rematerialization to improve results. Unlike deep ML compilers that organize memory allocation and deallocation at the computation graph level, STWeaver manages memory requests at the level of the overall model execution. These approaches are complementary and orthogonal with our work.
@@ 10. Conclusion
This work presents the design, implementation, and evaluation of STWeaver, a novel memory allocation system that significantly improves the memory utilization of large-scale model training. STWeaver builds on our insight of spatio-temporal regularity in model training allocation requests to combine ahead-of-time memory layout planning with runtime profile-guided allocation. Extensive evaluations show that STWeaver significantly outperforms state-of-the-art solutions in terms of both effectiveness and efficiency. Complementary to existing runtime defragmentation methods, we believe STWeaver demonstrates the powerful potential of fusing proactive pre-runtime planning with reactive runtime decision-making.

### v2

@@ 9.4. Performance Breakdown
To understand the performance contribution of the static and dynamic allocators in STAlloc, we evaluate the performance breakdown of STAlloc when training the Qwen1.5-MoE-A2.7B model with the same setting in §9.2. To this end, we sequentially disable the dynamic allocator reusing Static Allocation Plan (mentioned as STAlloc w/o reuse), and the static allocator (mentioned as Caching Allocator, which is the vanilla PyTorch caching Allocator), and measure the corresponding memory efficiency in the above cases.
Static Allocator. The results in Figure 13 indicate that STAlloc with only the Static Allocation Plan reduces fragmentation memory by 70.2% compared to PyTorch Caching Allocator. This reduction in fragmentation memory accounts for 91% of the total fragmentation memory reduction achieved by the complete STAlloc system relative to PyTorch. Static planning accounts for the predominant share of the defragmentation result, primarily because static memory allocations form a substantial majority (from 73.4% to 99.3%) of the total memory allocation, as shown in Table 3.
@@
Dynamic fallback w/o reuse (GB) 15.19 1.61 17.80 1.78 2.99 1.95
Dynamic fallback with reuse (GB) 15.19 1.12 17.22 1.70 1.92 1.55
Dynamic Allocator. Compared with STAlloc without dynamic reuse, the full STAlloc reduces memory fragmentation by an additional 22.9%, mainly by lowering fallback allocations to the caching allocator. As shown in Table 3, enabling dynamic reuse decreases the number of requests falling back to the caching allocator. This benefit is most evident under recomputation, where caching allocations drop by 24.9%. Without recomputation, the impact is smaller.
@@ 10. Related Work
Online GPU Allocators. To reduce memory fragmentation and improve allocation efficiency, a plethora of online GPU memory allocators (halloc; fast-dynamic-allocator; scatteralloc; register-alloc) have been developed. Dynamic allocators operate atop the native GPU memory APIs (e.g., cudaMalloc) in a similar manner as the caching allocator of PyTorch. Differently, such allocators are meant to run on GPU threads alongside GPGPU applications, rather than managing GPU memory from the host like PyTorch and GMLake. To reduce fragmentation, they usually adopt sophisticated allocation policies such as the slab and buddy systems. Also, to achieve high-throughput allocations on GPUs, they have proposed scalable synchronization primitives across the massive threads of GPUs (gelado). As a pluggable allocator of PyTorch, STAlloc also chooses to manage GPU memory from the host to improve usability and programmability.
Generic Memory Defragmentation Techniques. Memory defragmentation has been studied and discussed in various scenarios (fast-multicore-defrag; memory-frag-solved; eddy) beyond GPU applications. Previous work (parallel-dfrag-gpu; marlow-copy-based; sapphire-copy-based; siegwart-copy-based) has proposed defragmentation strategy based on data movement or copying. These approaches are mainly deployed in the real-time system with unpredictable runtime behaviors, which result in complex defragmentation strategies with high runtime overhead.
Machine Learning Compilers. Machine Learning compilers that convert high-level computation graphs to GPU instructions must manage memory allocation for these graphs. The compilers pre-analyze control dependencies to optimize memory layouts before execution. Compilers like TVM (tvm) and TFLite (tensorflow_micro) use greedy heuristics for reasonable allocation, while Checkmate (checkmate) employs solvers for optimal rematerialization to improve results. Unlike deep ML compilers that organize memory allocation and deallocation at the computation graph level, STAlloc manages memory requests at the level of the overall model execution. These approaches are complementary and orthogonal with our work.
@@ 11. Conclusion
This work presents the design, implementation, and evaluation of STAlloc, a novel memory allocation system that significantly improves the memory utilization of large-scale model training. STAlloc builds on our insight of spatio-temporal regularity in model training allocation requests to combine ahead-of-time memory layout planning with runtime profile-guided allocation. Extensive evaluations show that STAlloc significantly outperforms state-of-the-art solutions in terms of both effectiveness and efficiency. Complementary to existing runtime defragmentation methods, we believe STAlloc demonstrates the powerful potential of fusing proactive pre-runtime planning with reactive runtime decision-making.
Acknowledgements.
We thank the reviewers and our shepherd, Christoph Kirsch, for their insightful comments. This work was supported by the National Key R&D Program of China (2023YFB4502200), the National Natural Science Foundation of China (62325405, 62504139, U24B6015), Beijing Natural Science Foundation (QY24247, L242018, L257010), Beijing National Research Center for Information Science, Technology (BNRist), Beijing Innovation Center for Future Chips, and State Key laboratory of Space Network and Communications.
Appendix A Artifact
The artifact code is available at Zenodo:
https://zenodo.org/records/17173036